OpenStudy (anonymous):

Find the length of the radius of a sphere with a volume of 14.14 mm3.

OpenStudy (anonymous):

answers: 3.6 mm 1.8 mm 3.0 mm 1.5 mm

OpenStudy (blurbendy):

Volume of a sphere = 4/3(pi)r^3 You need to solve for r

OpenStudy (anonymous):

I don't know what r is tho and im so confuseddd.

OpenStudy (blurbendy):

r stands for for radius, which is what you're trying to find. to start, just substitute known values into the formula. 14.14 = 4/3(pi)r^3

OpenStudy (anonymous):

oh wait I do know what r is but I don't know what to do .

OpenStudy (blurbendy):

you'd start by dividing both sides by 4/3pi

OpenStudy (anonymous):

that gives me 3.7

OpenStudy (blurbendy):

i got something around 3.38

OpenStudy (blurbendy):

so you have 3.38 = r^3 now take the cubed root of both sides and you'll then have what the radius is

OpenStudy (anonymous):

1.8?

OpenStudy (blurbendy):

you can rewrite the cubed root of 3.38 as 3.38^(1/3)

OpenStudy (blurbendy):

which is... ?

OpenStudy (anonymous):

it gave me 1.1 ?

OpenStudy (anonymous):

ohhhh wow this is so confusing so itd be 1.5?

OpenStudy (blurbendy):

yes. just be careful when you're entering things in your calculator. Parentheses are important

OpenStudy (anonymous):

ohk thanks :)

OpenStudy (blurbendy):

yw